OpenStudy (anonymous):

Show all work to simplify (x/2)-(2/(x-1))+(2/(x-2)) and list the restrictions

hero (hero):

Hint: x = 1 and x = 2 will make the denominators zero

OpenStudy (anonymous):

is the common denominator x(x-2)?

hero (hero):

If x(x-2) was the common denominator, you would be able to use it to combine each fraction. But there are THREE fractions, and THREE different denominators, which means there will be THREE numbers multiplied together to form the LCM. x(x-2) represents only two numbers

OpenStudy (anonymous):

oh, then whats the lowest common denominator?

hero (hero):

In this case, LCM = LCD

OpenStudy (anonymous):

and what number is that? because i rele dont understand

hero (hero):

Hmmm. You have three different fractions, each with three different denominators, yet the denominators of each must be the same before you can combine them. What if you had \(\frac{1}{2} + \frac{2}{3} + \frac{3}{5}\)

hero (hero):

What would you do in this case?

OpenStudy (anonymous):

you multiply all of them but that would be 2x^2+3x+2

OpenStudy (anonymous):

you multiply all of them but that would be 2x^2+3x+2

hero (hero):

Yes, but guess what? The denominators would be the same and you would be able to combine them.

hero (hero):

Did you multiply 2(x-1)(x-2) ?

OpenStudy (anonymous):

i messed up with writing it it should be (2/x)-(2/x-1)+(2/x-2)

hero (hero):

No, you should write it like this 2/x - 2/(x-1) + 2/(x-2) to distinguish the numerator from the denominator.

hero (hero):

That becomes \(\large\frac{2}{x} - \frac{2}{x-1} + \frac{2}{x-2} \)

OpenStudy (anonymous):

im now really confused, what is the LCD

hero (hero):

The LCD is x(x-1)(x-2)